OpenStudy (mony01):

With what minimum speed must you toss a 160g ball straight up to just touch the 12-m-high roof of the gymnasium if you release the ball 1.4m above the ground? Solve this problem using energy.

OpenStudy (perl):

try to use the work energy theorem PEi + KEi = PEf + KEf

OpenStudy (perl):

KE= 1/2*mv^2 PE = m*g*h change 160 grams = .160 kg .160 * 9.8*1.4 + 1/2 * .160 * v^2 = .160 * 9.8 * (12) + 1/2*.160*0

OpenStudy (perl):

so energy is conserved

OpenStudy (perl):

1/2 * .160 * v^2 = .160 * 9.8 * (12) - .160 * 9.8*1.4

OpenStudy (perl):

i get v = 14.41 m/s

OpenStudy (perl):

do you have any questions about this . when the ball 'just touches' the ceiling, the velocity is zero